OpenStudy (anonymous):

simplify (3/5y^4)^-2

OpenStudy (luigi0210):

When you have a negative exponent you have to flip the problem to make it positive

OpenStudy (anonymous):

\[(\frac{3}{5y^4})^{-2}\] \[(\frac{5y^4}{3})^2\] Use the rule \[(\frac{m}{n})^p=\frac{m^p}{n^p}\] solve

OpenStudy (anonymous):

6/10y^16?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

10y^16/6?

OpenStudy (anonymous):

Well if you've used the rule, you would get \[(\frac{5y^4}{3})^2 \\ =\frac{5^2(y^4)}{3^2}^2\]

OpenStudy (anonymous):

That is not even an answer choice on my test, I asked Simplify!

OpenStudy (anonymous):

Yes, what is \[5^2\]

OpenStudy (anonymous):

25

OpenStudy (anonymous):

Good, what is \[(y^4)^2\]

OpenStudy (anonymous):

y8

OpenStudy (anonymous):

What about \[3^2\]?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

That became \[\frac{25y^8}{9}\]

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Best Responce

OpenStudy (anonymous):

:)